Formulate a differential equation whose general solution is z(t)=A sin t-B cos t.


1
Expert's answer
2021-04-29T14:51:24-0400

"Answer: z''+z=0"

"Check: \\
z'' + z = z''(t) + z(t) = (A \sin(t) - B\cos(t))'' + A\sin(t) - B\cos(t) = \\
= (A\cos(t) + B\sin(t))' + A\sin(t) - B\cos(t) = \\
= -A\sin(t) + B\cos(t) + A\sin(t) - B\cos(t) = 0"


If you decide strictly (optional):

"\begin{vmatrix}
 z(t) & z(t)' & z(t)'' \\
 sin(t) & cos(t) & -sin(t) \\
 -cos(t) & sin(t) & cos(t)
\end{vmatrix} = 0 \\
\Rightarrow z cos^2(t) + z'sin(t)cos(t) + z''sin^2(t) + z''cos^2(t) + zsin^2(t) - z'sin(t)cos(t) = 0\\
\Rightarrow z (cos^2(t) + sin^2(t)) + z''(sin^2(t) + cos^2(t)) = 0 \;\; \;\; (sin^2(t) + cos^2(t) = 1)\\
\Rightarrow z + z'' = 0"